# A Modular, Adaptive, and Scalable Quantum Factoring Algorithm

### v1

@@ 4 Computational Complexity
A key advantage of our proposed approach is a significant reduction in the number of phase (or counting) qubits per circuit block. For an $n$-bit integer, the size of the phase register in the standard algorithm is approximately $2n+1$, whereas in the modular framework it is reduced to $m_{\max}$, where $m_{\max}=\max(m_{1},m_{2},\ldots,m_{B})$ and $m_{k}$ is the number of qubits allocated to the $k$-th block. In practice, $m_{\max}$ can be chosen to be very small, for example 3 or 4 (ref. Section 3). At RSA-2048 scale ($n=2048$), the phase register requirement drops from 4097 to $m_{\max}$.
@@ 5 Conclusion
We have introduced a modular, windowed formulation of Shor’s algorithm that restructures phase estimation into $B$ independent, shallow blocks that can be executed sequentially or in parallel, followed by carry-aware classical stitching (ref. Algorithms 1, 2, 3 and 4). The result is a significant reduction of the counting qubit requirement (corresponding to the phase register) from approximately $2n+1$ to $m_{\max}$, where $m_{\max}=\max(m_{1},m_{2},\ldots,m_{B})$ and $m_{k}$ is the number of qubits allocated to the $k$-th block. Here, $m_{\max}$ can be chosen to be small, e.g. 3 or 4 (ref. Section 3). Further, these independent blocks are amenable to parallelization, which allows for a shallower effective circuit depth and significantly lower overall execution time, making the approach far more viable for near-term quantum hardware than the standard Shor’s algorithm. The number of qubits required for the work register in our approach remains the same as in the standard Shor’s algorithm.

### v5

@@ 6 Computational Complexity
[…]
A key advantage of our proposed approach is a significant reduction in the number of phase (or counting) qubits per circuit block. For an $n$-bit integer, the size of the phase register in the standard algorithm is approximately $2n+1$, whereas in the modular framework it is reduced to $m_{\max}$, where $m_{\max}=\max(m_{1},m_{2},\ldots,m_{B})$ and $m_{k}$ is the number of qubits allocated to the $k$-th block. In practice, $m_{\max}$ can be chosen to be very small, for example 3 or 4 phase qubits were sufficient for the computational examples considered (ref. Section 3). At RSA-2048 scale ($n=2048$), the phase register requirement drops from 4097 to $m_{\max}$.
@@ 7 Conclusion
We have introduced a modular, windowed formulation of Shor’s algorithm that restructures phase estimation into $B$ independent, shallow blocks that can be executed sequentially or in parallel, followed by carry-aware classical stitching (ref. Algorithms 1, 2, 3 and 4). The result is a significant reduction of the counting qubit requirement (corresponding to the phase register) from approximately $2n+1$ to $m_{\max}$, where $m_{\max}=\max(m_{1},m_{2},\ldots,m_{B})$ and $m_{k}$ is the number of qubits allocated to the $k$-th block. Here, $m_{\max}$ can be chosen to be small, e.g. 3 or 4 qubits were found to be sufficient for the computational examples considered (ref. Section 3). Further, these independent blocks are amenable to parallelization, which allows for a shallower effective circuit depth and significantly lower overall execution time, making the approach far more viable for near-term quantum hardware than the standard Shor’s algorithm. The number of qubits required for the work register in our approach remains the same as in the standard Shor’s algorithm.